OpenStudy (tkriz972):

Consider this system of linear equations: y − x = -13 -4x + 3y = -51 The solution of the system of equations is

OpenStudy (anonymous):

y=x-13. Plug this into the bottom equation.

OpenStudy (anonymous):

-4x+3(x-13)=-51

OpenStudy (tkriz972):

those arnt awnser choices

OpenStudy (anonymous):

y = x-13 from the first equation Replace y in the second equation with x-13 and then solve the result for x.

OpenStudy (tkriz972):

ight appreciate it

OpenStudy (anonymous):

lol i pretty much made the equation for you, u just gotta multply and add the x's, then divide from 51 :)

OpenStudy (tkriz972):

aye i got the right awnser thx

OpenStudy (tkriz972):

oh my bad im not very clever lol

OpenStudy (anonymous):

Np. Main objective is to find the y or x of one of them (basically y= whatever, or x=whatever). you plug whatever in the other equation